OpenStudy (anonymous):

What is the least common multiple (LCM) of 25mn2 , 15mn, and 10m2n2? A. 5mn B. 5m2 :D n 2 C. 150mn D. 150m2 n

OpenStudy (gorv):

mn is common in all three term agree??

OpenStudy (anonymous):

yeah then

OpenStudy (gorv):

now 25 15 and 10 5 is common in them

OpenStudy (gorv):

if take out 5

OpenStudy (gorv):

5 , 3 ,2

OpenStudy (gorv):

their lcm= 5*3*2=30

OpenStudy (anonymous):

yeah I know that but what about mn

OpenStudy (gorv):

after takin out mn we left with that is 5mn

OpenStudy (gorv):

5mn{5n,3,2mn}

OpenStudy (gorv):

so their lcm=5mn*(5n*3*2mn)

OpenStudy (anonymous):

okay got iit

OpenStudy (gorv):

5*3*2 we calculated.....and mn*n=mn^2 so 5mn*30mn^2=150m^2*n^3

OpenStudy (anonymous):

thankyou :D

OpenStudy (gorv):

welcome:)